# Traversal Verification for Speculative Tree Decoding

### v1

@@ 4.2 Overall Effectiveness
We present the acceptance lengths and throughput of two combinations of draft and target model, namely Llama3.2-1B-Instruct with Llama3.1-8B-Instruct and Llama-68M with Llama2-7B in Table 2 and Table 3. For chain and binary tree, we set the depth at 5, which is equal to the maximum depth of EAGLE sparse tree. Tok.V denotes token-level verification and Tra.V denotes Traversal Verification. The acceptance lengths are rounded to 2 decimal places, and we also provide the standard errors. $\Delta$ denotes the relative improvement of Traversal Verification over token-level verification. The baseline generation speed without speculative decoding for Llama3.1-8B-Instruct is 34.5$\pm$0.1 token/s and for Llama2-7B is 37.3$\pm$0.1 token/s, and the speedup ratio can be calculated accordingly.
@@
Multi-turn 3.95$\pm$0.03 4.09$\pm$0.03 3.5% 4.64$\pm$0.05 4.76$\pm$0.04 2.6% 4.53$\pm$0.02 4.67$\pm$0.02 3.1%
Translation 3.50$\pm$0.02 3.53$\pm$0.04 1.0% 4.28$\pm$0.02 4.43$\pm$0.03 3.4% 4.16$\pm$0.04 4.27$\pm$0.03 2.6%
Sum. 3.66$\pm$0.02 3.76$\pm$0.03 2.6% 4.51$\pm$0.02 4.64$\pm$0.02 2.7% 4.32$\pm$0.03 4.46$\pm$0.03 3.1%
QA 3.51$\pm$0.02 3.68$\pm$0.03 4.7% 4.32$\pm$0.05 4.40$\pm$0.04 2.0% 4.19$\pm$0.05 4.31$\pm$0.06 2.9%
Math 4.61$\pm$0.05 4.70$\pm$0.03 1.8% 5.37$\pm$0.03 5.39$\pm$0.05 0.4% 5.13$\pm$0.01 5.21$\pm$0.02 1.5%
RAG 4.05$\pm$0.04 4.17$\pm$0.05 3.1% 4.63$\pm$0.02 4.76$\pm$0.06 2.8% 4.60$\pm$0.03 4.68$\pm$0.04 1.7%
Avg. Accept. 3.88$\pm$0.02 3.99$\pm$0.01 2.8% 4.63$\pm$0.03 4.73$\pm$0.01 2.2% 4.49$\pm$0.02 4.60$\pm$0.02 2.4%
Avg. Token/s 51.2$\pm$1.2 52.5$\pm$1.1 2.5% 54.0$\pm$0.6 54.9$\pm$1.2 1.7% 57.3$\pm$1.3 58.5$\pm$0.8 2.1%
@@
Multi-turn 2.05$\pm$0.05 2.16$\pm$0.03 5.5% 2.47$\pm$0.01 2.59$\pm$0.01 4.7% 2.55$\pm$0.02 2.70$\pm$0.02 5.6%
Translation 1.97$\pm$0.05 2.10$\pm$0.05 6.3% 2.38$\pm$0.01 2.43$\pm$0.03 2.1% 2.49$\pm$0.01 2.51$\pm$0.03 0.9%
Sum. 1.77$\pm$0.04 1.86$\pm$0.05 4.9% 2.14$\pm$0.01 2.27$\pm$0.03 5.8% 2.25$\pm$0.02 2.36$\pm$0.02 4.7%
QA 2.07$\pm$0.01 2.19$\pm$0.02 5.6% 2.59$\pm$0.05 2.71$\pm$0.01 4.8% 2.63$\pm$0.02 2.69$\pm$0.02 2.2%
Math 2.01$\pm$0.05 2.15$\pm$0.04 7.0% 2.49$\pm$0.05 2.67$\pm$0.06 7.0% 2.57$\pm$0.02 2.72$\pm$0.01 6.0%
RAG 2.09$\pm$0.05 2.19$\pm$0.03 4.8% 2.56$\pm$0.05 2.69$\pm$0.05 5.0% 2.63$\pm$0.02 2.71$\pm$0.06 3.2%
Avg. Accept. 1.99$\pm$0.01 2.10$\pm$0.01 5.7% 2.44$\pm$0.03 2.56$\pm$0.01 4.9% 2.52$\pm$0.01 2.62$\pm$0.01 3.8%
Avg. Token/s 58.0$\pm$0.7 60.8$\pm$0.8 4.8% 59.4$\pm$0.8 61.6$\pm$0.6 3.7% 69.1$\pm$0.9 71.2$\pm$1.0 3.0%
@@ 4.3 Impact of Chain Depth and Tree Size
Since Traversal Verification considers the joint probability of the entire sequence, its performance improvement becomes more pronounced as the tree size and depth increases. Specifically, for chain decoding, we conduct experiments at depths of 2, 4, 6, and 8. For tree decoding, we employ binary trees from depths of 2 to 5 (corresponding to trees with $2^{3}$-1, $2^{4}$-1, $2^{5}$-1, and $2^{6}$-1 nodes, respectively).
@@
As shown in Figure 3, the advantage of Traversal Verification grows progressively with increasing chain depth and tree size. In specialized scenarios (e.g., model offloading) where large tree sizes are permissible (for example, Sequoia [4] utilizes trees with 768 nodes and depth exceeding 20), Traversal Verification is expected to demonstrate even greater performance gains.
@@ 5 Related work
Significant efforts have been devoted to accelerating LLMs. Some approaches directly reduce memory access and computational costs through techniques such as quantization [7, 8, 32, 21] and knowledge distillation [10, 16, 36]. Some other works focus on architectural innovations, such as Mixture of Experts (MoE) [14, 6], where only a subset of model parameters is activated during inference, thereby improving inference speed.
Speculative decoding [3, 18] introduces a distinct drafting-verification paradigm that leaves the LLM itself unchanged. Researches on speculative decoding primarily focus on two directions. 1) Better alignment between the draft and the target model, such as EAGLE [20, 19] and Medusa [2] series. 2) Better verification strategies, such as innovations in tree structures [19, 4, 30] and verification algorithms, which are more closely related to this work.
In chain decoding scenarios, Block Verification [26] and Asps [11] identify the sub-optimality in token-level verification and propose enhancements. SpecTr [27] extends chain decoding to multi-candidate settings by formulating it as an optimal transport problem solved via linear programming, while SpecInfer [22] employs Recursive Rejection Sampling for multi-candidate situations. Subsequent works refine this approach into RRSw (recursive rejection sampling without replacement) [4, 20, 13, 34], preventing repeated sampling and rejection of identical tokens, thereby improving acceptance rates. Beyond standard sampling, SpecHub [25] and Greedy Sampling [12] adopt hybrid strategies: deterministically selecting top-K candidates with the highest probability and sampling other candidates probabilistically, achieving higher acceptance rates in specific scenarios.
@@ Appendix A Formal Proof of Losslessness of Traversal Verification
We first prove a necessary and sufficient condition for a valid tree verification algorithm (Definition 3.2). Our proof technique is analogous to [26, Lemma 2 in Appendix.B] and extends the original lemma to a tree-structured case.
@@
$\forall\mathcal{M}_{s},\mathcal{M}_{b}$, let $T$ be a valid sampling tree on $\mathcal{M}_{s}$ with root node $X_{0}$, and $\gamma_{\max}:=\max_{v\subset T}\gamma_{v}$ be the maximum depth along all chains in $T$. The output of a tree verification algorithm $\mathcal{A}_{\mathrm{ver}}$ is denoted as
@@
with $\widetilde{Z}_{>\widetilde{\tau}+1}\sim\mathcal{M}_{b}(\cdot|\widetilde{X}^{\widetilde{\tau}},\widetilde{Y})$. Note that by Equation (5), we have
@@
$\forall\mathcal{M}_{s},\mathcal{M}_{b}$, let $T$ be a valid sampling tree on $\mathcal{M}_{s}$ with root node $X_{0}$ and $\gamma_{\max}:=\max_{v\subset T}\gamma_{v}$ be the maximum depth along all chains in $T$. The first chain in $T$ is denoted as $\alpha=(\alpha_{0},\alpha_{1},\dots,\alpha_{\gamma_{\alpha}})$ from root $\alpha_{0}=X_{0}$ to the first leaf node $\alpha_{\gamma_{\alpha}}$. $Z^{\gamma_{\max}}$ is the sequence generated by Traversal Verification $\mathcal{A}_{\rm tra}(T,\mathcal{M}_{s},\mathcal{M}_{b})$ (i.e., Algorithm 3) in Lemma A.1. Then the following statements hold, $\forall 0\leqslant\ell\leqslant\gamma_{\alpha}$,
@@
$\mathbb{P}(Z^{\gamma_{\max}}=(\alpha^{\ell},Z_{>\ell})|Z_{\ell+1}\neq\alpha_{\ell+1})$ $=\mathbb{P}(Z^{\gamma_{\max}}|Z^{\ell}=\alpha^{\ell}\text{ in }T_{\rm new})$
@@
$\mathbb{P}(Z^{\ell}=\alpha^{\ell})$ =(9)${=}p^{ini}_{\alpha}(\alpha_{\ell+1})+(1-p^{ini}_{\alpha}(\alpha_{\ell+1}))\cdot\mathbb{P}(Z^{\ell}=\alpha^{\ell}|Z_{\ell+1}\neq\alpha_{\ell+1})$ )
=(13)${=}p^{ini}_{\alpha}(\alpha_{\ell+1})+(1-p^{ini}_{\alpha}(\alpha_{\ell+1}))\cdot p^{\prime}_{\alpha}(\alpha_{\ell}).$ ) . (15)
@@
[…]
${=}$ $\sum_{x}\min\left\{p^{ini}_{\alpha}(\alpha_{\ell})\mathcal{M}_{b}(x|\alpha^{\ell}),\mathcal{M}_{s}(x|\alpha^{\ell})\right\}+\sum_{x}\left[p^{ini}_{\alpha}(\alpha_{\ell})\mathcal{M}_{b}(x|\alpha^{\ell})-\mathcal{M}_{s}(x|\alpha^{\ell})\right]_{+}$
[…]
@@
${=}$ $\mathcal{M}_{s}(x_{\ell+1}|\alpha^{\ell})\cdot p^{ini}_{\alpha}(\alpha_{\ell+1}=x_{\ell+1})\mathcal{M}_{b}(x_{>{\ell+1}}|\alpha^{\ell},x_{\ell+1})$
$+\sum_{x\neq x_{\ell+1}}\mathcal{M}_{s}(x^{\prime}|\alpha^{\ell})[1-p^{ini}_{\alpha}(\alpha_{\ell+1}=x^{\prime})]\mathbb{P}(Z^{\gamma_{\max}}=(\alpha^{\ell},x_{>\ell})|x_{\ell+1}\neq\alpha_{\ell+1})$
${=}$ $\mathcal{M}_{s}(x_{\ell+1}|\alpha^{\ell})\cdot p^{ini}_{\alpha}(\alpha_{\ell+1}=x_{\ell+1})\mathcal{M}_{b}(x_{>{\ell+1}}|\alpha^{\ell},x_{\ell+1})$
@@
[…]
@@
    (A.1)$=\mathcal{M}_{b}(x_{>{\ell+1}}|\alpha^{\ell},x_{\ell+1})\cdot p^{ini}_{\alpha}(\alpha_{\ell})\mathcal{M}_{b}(x_{\ell+1}|\alpha^{\ell})=p^{ini}_{\alpha}(\alpha_{\ell})\mathcal{M}_{b}(x_{>{\ell}}|\alpha^{\ell}).$ ) .
@@
    Since $[\mathcal{M}_{s}(x_{\ell+1}|\alpha^{\ell})-p^{ini}_{\alpha}(\alpha_{\ell})\mathcal{M}_{b}(x_{\ell+1}|\alpha^{\ell})]_{+}=0$, we have
@@
[…]
    Taking it into (*A.1), then
[…]
In conclusion, (A.1)$=p^{ini}_{\alpha}(\alpha_{\ell})\mathcal{M}_{b}(x_{>{\ell}}|\alpha^{\ell})$ ), i.e.,
@@
Appendix B Traversal Verification for Speculative Chain Decoding
B.1 Formal Proof of Theorem 3.4
To establish the optimality of Traversal Verification in the single-chain case, we need to introduce two lemmas presented in [26].
Lemma B.1 (Lemma 3 in [26]).
Let $T=(\alpha_{0},\dots,\alpha_{\gamma})$ be a stochastic sampling chain based on $\mathcal{M}_{s}$, $\mathcal{A}_{block}$ be Block Verification proposed in [26], and
@@ Appendix B Traversal Verification for Speculative Chain Decoding
Then we have $\forall\ell\leqslant\gamma$, and sequence $x^{\ell}$,
$\mathbb{P}(\tau\geqslant\ell|\alpha^{\ell}=x^{\ell})=p_{x}^{ini}(x_{\ell}),$
which means that
$\mathbb{P}(\tau\geqslant\ell|\alpha^{\ell})=p_{\alpha}^{ini}(\alpha_{\ell}).$
Lemma B.2 (Lemma 4 in [26]).
For chain verification algorithms that satisfy the constraints in Lemma A.2, we have $\forall\ell\leqslant\gamma$, and sequence $x^{\ell}$,
$\mathbb{P}(\tau\geqslant\ell|\alpha^{\ell}=x^{\ell})\leqslant p_{x}^{ini}(x_{\ell}),$
which means that
$\mathbb{P}(\tau\geqslant\ell|\alpha^{\ell})\leqslant p_{\alpha}^{ini}(\alpha_{\ell}).$
@@
It suffices to observe that when the stochastic sampling tree reduces to a single-chain structure, the equivalent definition Lemma A.1 of valid chain verification algorithm in this paper is identical to the equivalent definition [26, Lemma 2] of the valid draft verification algorithm. Therefore, Lemmas B.1 and B.2 hold automatically as the directly applications of [26, Lemma 3 and Lemma 4]. ∎
@@
$\mathbb{P}(Z^{\ell}=\alpha^{\ell})=p^{ini}_{\alpha}(\alpha_{\ell}),$
then we have
$\mathbb{P}(\tau\geqslant\ell|\alpha^{\ell})=\mathbb{P}(Z^{\ell}=\alpha^{\ell})=p^{ini}_{\alpha}(\alpha_{\ell}).$
By Lemmas B.1 and B.2, we show that among all valid chain verification algorithms (i.e., valid draft verification algorithms satisfying the constraints in [26, Lemma 2]), Traversal Verification accepts any given subsequence with the highest probability as the same as Block Verification. Specifically, for any given sampling sequence $\alpha$, we have
$\mathbb{E}[N_{\rm traversal}]$ $=\sum_{\ell}\mathbb{P}(\tau_{\rm traversal}\geqslant\ell|\alpha^{\ell})=\sum_{\ell}p_{\alpha}^{ini}(\alpha_{\ell}),$
$\mathbb{E}[N_{\rm block}]$ $=\sum_{\ell}\mathbb{P}(\tau_{\rm block}\geqslant\ell|\alpha^{\ell})=\sum_{\ell}p_{\alpha}^{ini}(\alpha_{\ell}),$
$\mathbb{E}[N_{\rm verify}]$ $=\sum_{\ell}\mathbb{P}(\tau_{\rm verify}\geqslant\ell|\alpha^{\ell})\leqslant\sum_{\ell}p_{\alpha}^{ini}(\alpha_{\ell}).$

### v2

@@ 4.2 Overall Effectiveness
[…]
@@
[…]
@@
[…]
@@ 4.3 Impact of Chain Depth and Tree Size
Since Traversal Verification considers the joint probability of the entire sequence, it is intuitive that the performance improvement will become more pronounced as the tree size and depth increase. To illustrate these effects, we perform experiments across varying chain depths and tree sizes. Specifically, for chain decoding, we conduct experiments at depths of 2, 4, 6, and 8. For tree decoding, we employ binary trees from depths of 2 to 5 (corresponding to trees with $2^{3}$-1, $2^{4}$-1, $2^{5}$-1, and $2^{6}$-1 nodes, respectively).
@@
As shown in Figure 3, the advantage of Traversal Verification grows progressively with increasing chain depth and tree size. In specialized scenarios (e.g., model offloading) where large tree sizes are permissible (for example, Sequoia [ChenSequoia] utilizes trees with 768 nodes and depth exceeding 20), Traversal Verification is expected to demonstrate even greater performance gains.
@@ 5 Related work
Significant efforts have been devoted to accelerating LLMs. Some approaches directly reduce memory access and computational costs through techniques such as quantization [dettmers2022llmint88bitmatrixmultiplication, frantar2023gptqaccurateposttrainingquantization, DBLP:conf/icml/XiaoLSWDH23, DBLP:conf/mlsys/0002TTYCWXDG024] and knowledge distillation [gu2024minillm, DBLP:conf/icml/KoKCY24, DBLP:conf/acl/Zhong00L0T24]. Some other works focus on architectural innovations, such as Mixture of Experts (MoE) [mixtral, dsv3], where only a subset of model parameters is activated during inference, thereby improving inference speed.
Speculative decoding [DBLP:journals/corr/abs-2302-01318, DBLP:conf/icml/LeviathanKM23] introduces a distinct drafting-verification paradigm that leaves the LLM itself unchanged. Researches on speculative decoding primarily focus on two directions. 1) Better alignment between the draft and the target model, such as EAGLE [LiEAGLE, LiEAGLE2] and Medusa [Medusa] series. 2) Better verification strategies, such as innovations in tree structures [LiEAGLE2, ChenSequoia, optree] and verification algorithms, which are more closely related to this work.
In chain decoding scenarios, Block Verification [blockV] and Asps [Asps] identify the sub-optimality in token-level verification and propose enhancements. SpecTr [SpecTr] extends chain decoding to multi-candidate settings by formulating it as an optimal transport problem solved via linear programming, while SpecInfer [DBLP:conf/asplos/MiaoOZCWZWZYSSC24] employs Recursive Rejection Sampling for multi-candidate situations. Subsequent works refine this approach into RRSw (recursive rejection sampling without replacement) [ChenSequoia, LiEAGLE, JeonRSD, YangMCSD], preventing repeated sampling and rejection of identical tokens, thereby improving acceptance rates. Beyond standard sampling, SpecHub [SpecHub] and Greedy Sampling [TowardsOptimal] adopt hybrid strategies: deterministically selecting top-K candidates with the highest probability and sampling other candidates probabilistically, achieving higher acceptance rates in specific scenarios.
@@ 6 Conclusion
Acknowledgments and Disclosure of Funding
This project is fully funded by Lenovo. We would like to express special thanks to the Lenovo AI Lab and the Lenovo Model Factory Team for their valuable support in providing computing resources.
@@ Appendix A Formal Proof of Losslessness of Traversal Verification
We first prove a necessary and sufficient condition for a valid tree verification algorithm (Definition 3.2). Our proof technique is analogous to [blockV, Lemma 2 in Appendix.B] and extends the original lemma to a tree-structured case.
@@
$\forall\mathcal{M}_{s},\mathcal{M}_{b}$, let $T$ be a decoding tree rooted at $X_{0}$ base on $\mathcal{M}_{s}$, and $\gamma_{\max}:=\max_{v\subset T}\gamma_{v}$ be the maximum depth along all chains in $T$. The output of a tree verification algorithm $\mathcal{A}_{\mathrm{ver}}$ is denoted as
@@
[…]
@@
$\forall\mathcal{M}_{s},\mathcal{M}_{b}$, let $T$ be a decoding tree rooted at $X_{0}$ base on $\mathcal{M}_{s}$ and $\gamma_{\max}:=\max_{v\subset T}\gamma_{v}$ be the maximum depth along all chains in $T$. The first chain in $T$ is denoted as $\alpha=(\alpha_{0},\alpha_{1},\dots,\alpha_{\gamma_{\alpha}})$ from root $\alpha_{0}=X_{0}$ to the first leaf node $\alpha_{\gamma_{\alpha}}$. $Z^{\gamma_{\max}}$ is the sequence generated by Traversal Verification $\mathcal{A}_{\rm tra}(T,\mathcal{M}_{s},\mathcal{M}_{b})$ (i.e., Algorithm 3) in Lemma A.1. Then the following statements hold, $\forall 0\leqslant\ell\leqslant\gamma_{\alpha}$,
@@
$\mathbb{P}(Z^{\gamma_{\max}}=(\alpha^{\ell},Z_{>\ell})|Z_{\ell+1}\neq\alpha_{\ell+1})$ $=\mathbb{P}(Z^{\gamma_{\max}}=(\alpha^{\ell},Z_{>\ell})|T_{\rm new})$
@@
[…]
@@
$=$~ $\mathbb{E}_{\alpha_{\ell+1}}[\mathbb{P}(Z^{\ell}=\alpha^{\ell})]$
$=$~ $\sum_{x}p^{ini}_{\alpha}(\alpha_{\ell+1}=x)\mathcal{M}_{s}(x|\alpha^{\ell})+\left[1-\sum_{x}p^{ini}_{\alpha}(\alpha_{\ell+1}=x)\mathcal{M}_{s}(x|\alpha^{\ell})\right]\cdot p^{\prime}_{\alpha}(\alpha_{\ell})$
$=$~ $\sum_{x}\min\left\{p^{ini}_{\alpha}(\alpha_{\ell})\mathcal{M}_{b}(x|\alpha^{\ell}),\mathcal{M}_{s}(x|\alpha^{\ell})\right\}+p^{\prime}_{\alpha}(\alpha_{\ell})\sum_{x}\left[\mathcal{M}_{s}(x|\alpha^{\ell})-p^{ini}_{\alpha}(\alpha_{\ell})\mathcal{M}_{b}(x|\alpha^{\ell})\right]_{+}$
@@
$=$~ $\sum_{x}\left[p^{ini}_{\alpha}(\alpha_{\ell})\mathcal{M}_{b}(x|\alpha^{\ell})+\mathcal{M}_{s}(x|\alpha^{\ell})\right]-\sum_{x}\mathcal{M}_{s}(x|\alpha^{\ell})$
$=$~ $p^{ini}_{\alpha}(\alpha_{\ell}).$
@@
$=$~ $\mathbb{P}(\alpha_{\ell+1}=x_{\ell+1})\mathbb{P}(Z^{\gamma_{\max}}=(\alpha^{\ell+1},x_{>\ell+1})|\alpha_{\ell+1}=x_{\ell+1})$
@@
$+\sum_{x^{\prime}\neq x_{\ell+1}}\mathcal{M}_{s}(x^{\prime}|\alpha^{\ell})[1-p^{ini}_{\alpha}(\alpha_{\ell+1}=x^{\prime})]\mathbb{P}(Z^{\gamma_{\max}}=(\alpha^{\ell},x_{>\ell})|x_{\ell+1}\neq\alpha_{\ell+1})$
[…]
@@
$=$~ $\mathcal{M}_{b}(x_{>{\ell+1}}|\alpha^{\ell},x_{\ell+1})\cdot\min\left\{p^{ini}_{\alpha}(\alpha_{\ell})\mathcal{M}_{b}(x_{\ell+1}|\alpha^{\ell}),\mathcal{M}_{s}(x_{\ell+1}|\alpha^{\ell})\right\}$
@@
[…]
@@
    Since $\mathcal{M}_{s}(x_{\ell+1}|\alpha^{\ell})-p^{ini}_{\alpha}(\alpha_{\ell})\mathcal{M}_{b}(x_{\ell+1}|\alpha^{\ell})]_{+}=0$, we have
@@
    $=$~ $p^{\prime}_{\alpha}(\alpha_{\ell})\mathcal{M}^{\prime}_{b}(x_{\ell+1}|\alpha^{\ell})\left\{\sum_{x^{\prime}}\left[p^{ini}_{\alpha}(\alpha_{\ell})\mathcal{M}_{b}(x^{\prime}|\alpha^{\ell})-\mathcal{M}_{s}(x^{\prime}|\alpha^{\ell})\right]_{+}+1-p^{ini}_{\alpha}(\alpha_{\ell})\right\}$
    $=$~ $[p^{ini}_{\alpha}(\alpha_{\ell})\mathcal{M}_{b}(x_{\ell+1}|\alpha^{\ell})-\mathcal{M}_{s}(x_{\ell+1}|\alpha^{\ell})]_{+}$
    $=$~ $p^{ini}_{\alpha}(\alpha_{\ell})\mathcal{M}_{b}(x_{\ell+1}|\alpha^{\ell})-\mathcal{M}_{s}(x_{\ell+1}|\alpha^{\ell}).$
    Taking it into (*A.1), then
    (*A.1) $=\mathcal{M}_{b}(x_{>\ell+1}|\alpha^{\ell},x_{\ell+1})\left(\mathcal{M}_{s}(x_{\ell+1}|\alpha^{\ell})+p^{ini}_{\alpha}(\alpha_{\ell})\mathcal{M}_{b}(x_{\ell+1}|\alpha^{\ell})-\mathcal{M}_{s}(x_{\ell+1}|\alpha^{\ell})\right)$
@@
Appendix B Formal Proof of Single-chain Optimality
To establish the optimality of Traversal Verification in the single-chain case, we need to introduce two lemmas presented in [blockV].
Lemma B.1 (Lemma 3 in [blockV]).
Let $T=(\alpha_{0},\dots,\alpha_{\gamma})$ be a decoding chain based on $\mathcal{M}_{s}$, $\mathcal{A}_{block}$ be Block Verification proposed in [blockV], and
@@ Appendix B Formal Proof of Single-chain Optimality
Then we have $\forall\ell\leqslant\gamma$,
$\mathbb{P}(\tau\geqslant\ell|X^{\ell}=\alpha^{\ell})=p_{\alpha}^{ini}(\alpha_{\ell}).$
Lemma B.2 (Lemma 4 in [blockV]).
For chain verification algorithms that satisfy the constraints in Lemma A.2, we have $\forall\ell\leqslant\gamma$,
$\mathbb{P}(\tau\geqslant\ell|X^{\ell}=\alpha^{\ell})\leqslant p_{\alpha}^{ini}(\alpha_{\ell}).$
@@
It suffices to observe that when the stochastic sampling tree reduces to a single-chain structure, the equivalent definition Lemma A.1 of valid chain verification algorithm in this paper is identical to the equivalent definition [blockV, Lemma 2] of the valid draft verification algorithm. Therefore, Lemmas B.1 and B.2 hold automatically as the directly applications of [blockV, Lemma 3 and Lemma 4]. ∎
@@
$\mathbb{P}(\tau\geqslant\ell|X^{\ell}=\alpha^{\ell})=\mathbb{P}(Z^{\ell}=\alpha^{\ell})=p^{ini}_{\alpha}(\alpha_{\ell}).$
By Lemmas B.1 and B.2, we show that among all valid chain verification algorithms (i.e., valid draft verification algorithms satisfying the constraints in [blockV, Lemma 2]), Traversal Verification accepts any given subsequence with the highest probability as the same as Block Verification. Specifically, for a given decoding chain $\alpha=(\alpha_{0},\dots,\alpha_{\gamma})$ based on $\mathcal{M}_{s}$, we have
$\mathbb{E}[N_{\rm traversal}]$ $=\mathbb{E}_{\alpha\sim\mathcal{M}_{s}}\left[\sum_{\ell}\mathbb{P}(\tau_{\rm traversal}\geqslant\ell|X^{\ell}=\alpha^{\ell})\right]=\mathbb{E}_{\alpha\sim\mathcal{M}_{s}}\left[\sum_{\ell}p_{\alpha}^{ini}(\alpha_{\ell})\right],$
$\mathbb{E}[N_{\rm block}]$ $=\mathbb{E}_{\alpha\sim\mathcal{M}_{s}}\left[\sum_{\ell}\mathbb{P}(\tau_{\rm block}\geqslant\ell|X^{\ell}=\alpha^{\ell})\right]=\mathbb{E}_{\alpha\sim\mathcal{M}_{s}}\left[\sum_{\ell}p_{\alpha}^{ini}(\alpha_{\ell})\right],$
$\mathbb{E}[N_{\rm verify}]$ $=\mathbb{E}_{\alpha\sim\mathcal{M}_{s}}\left[\sum_{\ell}\mathbb{P}(\tau_{\rm verify}\geqslant\ell|X^{\ell}=\alpha^{\ell})\right]\leqslant\mathbb{E}_{\alpha\sim\mathcal{M}_{s}}\left[\sum_{\ell}p_{\alpha}^{ini}(\alpha_{\ell})\right].$